The admission fee at an amusement park is $1.25 for children and $7.20 for adults. On a certain day, 320 people entered the park, and the admission fees collected totaled 1114 dollars. How many children and how many adults were admitted?
Found 2 solutions by ikleyn, Boreal:
Answer by ikleyn(52756)

.

The TWIN problem was solved for you 5 minutes ago at this forum under the link

https://www.algebra.com/algebra/homework/word/finance/Money_Word_Problems.faq.question.1149478.html

https://www.algebra.com/algebra/homework/word/finance/Money_Word_Problems.faq.question.1149478.html


Use it as your TEMPLATE and solve the given problem by the same way.



Answer by Boreal(15235)

A+C=310 count equation and A=320-C
7.2A+1.25S=1114 money equation
7.2(320-C)+1.25C=1114
2304-7.2C+1.25C=1114
-5.95 C=-1190
C=200@1.25/C=$250
A=320-200=120 @7.20/A=$864
200 children and 120 adults